its dividing polynomials if its possible work would be nice
(6t^3+5t^2+9)/(2t+3)

Answer by stanbon(75887)

(6t^3+5t^2+9)/(2t+3)
----
(2t+3))6t^3+5t^2+9
-------
1st piece of the quotient: 3t^2
---
Multiply: 3t^2(2t+3) = 6t^3+9t^2
---
Subtract from 6t^3+5t^2+9 to get: -4t^2+9
---
2nd piece of the quotient: -2t
--------
Multiply: -2t(2t+3) = -4t^2-6t
-----
Subtract from 4t^2+9 to get: -6t-9
----
3rd piece of the quotient: -3
-------
Multiply -3(2t+3) = -6t-9
-------
Subtract from -6t-9 to get: 0
===========================
Final Answer: 3t^2-2t-3
===============================
Cheers,
Stan H.
=============
----